1. Find f(x) f(x) = 27 {rewrite first-leave answer as a radical) 3 +1 2. Find f() 2. Find f(x) f(x) = x {rewrite first; 3.
Answer #1

1. If f(x) = 2 (x343 get : differentiate with respect to x, f(x) = 2. (x)2/3-1 (x)/ 4 3 x 2. f(x) = +1 X4 3. x4 +1 we gett ff(x) = 2 . a la en(t). d (1+3x² + 3 det. d (4+32²) da dn 62 (6x) - 2. +3. et. Put the values of t and to + we get f(x) = += (2x+1) a (x²) - x h (2x+1) (2x+32 11 (2x+1) (2x) - x² (2) 2x 4 x² 420 – 2x² (2x+1)2 2x² +22 (2x+12 f(x) = an(n+1) (2x+1)2